# Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning

### v2

@@ B.8 Failure Cases
We demonstrate some failure cases of our method in Fig. E to show the boundary of the capability. We find it is challenging for our method to synthesize smooth camera viewpoint change in output videos (first example). If there is a big difference in viewpoints of two condition images, our method is very likely to treat the two images as irrelevant images and generates a hard transition without a reasonable motion. In addition, we also observe watermarks in a few output videos (second example). This is an inherent bias of the foundation model, which is inherited by our method. Hence, the performance of our method is still limited by the pretrained T2V foundation model.
C Implementation Details
Class Prompt Class Prompt
ApplyEyeMakeup “A person is applying eye makeup” ApplyLipstick “A person is applying lip stick”
BabyCrawling “A baby is crawling” BenchPress “A person is doing a bench press exercise”
Biking “A person is riding a bicycle” BlowDryHair “A person is blow drying their hair”
BodyWeightSquats “A person is performing bodyweight squats” BoxingPunchingBag “A person is punching a boxing bag”
BrushingTeeth “A person is brushing their teeth” CuttingInKitchen “A person is cutting food in the kitchen”
Drumming “A person is playing drums” Fencing “Two people are participating in fencing”
Haircut “A person is giving someone a haircut” HandstandPushups “A person is performing handstand pushups”
HeadMassage “A person is doing a head massage for another one” HorseRace “People are participating in a horse race”
HorseRiding “A person is riding a horse” JumpRope “A person is jumping rope”
Kayaking “A person is kayaking in water” Knitting “A person is knitting”
Mixing “A person is mixing ingredients” MoppingFloor “A person is mopping a floor”
PlayingCello “A person is playing a cello” PlayingFlute “A person is playing a flute”
PlayingGuitar “A person is playing a guitar” PlayingPiano “A person is playing a piano”
PlayingViolin “A person is playing a violin” PullUps “A person is doing pull-ups”
PushUps “A person is performing push-ups” Rafting “People are rafting in a river”
RockClimbingIndoor “A person is climbing an indoor rock wall” Rowing “People are rowing a boat”
ShavingBeard “A person is shaving their beard” SkateBoarding “A person is skateboarding”
Skiing “A person is skiing down a slope” Skijet “A person is riding a jet ski”
SkyDiving “A person is skydiving” Surfing “A person is surfing on a wave”
TrampolineJumping “A person is jumping on a trampoline” Typing “A person is typing on a keyboard”
WalkingWithDog “A person is walking a dog” WritingOnBoard “A person is writing on a board”
@@ C Implementation Details
Table E: Prompt for each selected action class in UCF-101. More details of the test set is elaborated in Sec. C.1.
@@ C.1 Establishment of Test Set
ni2024ti2v select 10 classes in UCF-101 which are not diverse enough for a reliable evaluation. The textual prompts used in their work are also oversimplified. In this work, we further extend the test set to 42 classes with manually annotated textual prompt for each action. For each class, there are videos of 25 subjects performing the same action. For diversity, we randomly select one video of each subject in each action class, totaling 1050 test data samples. The prompts are listed in Tab. E.
@@ C.3 Implementation of Previous Methods
D Limitation and Future Work
Though our approach achieves great performance in many experimental settings compared with previous methods, we still observe some limitations as illustrated in Fig. E. First of all, FlexTI2V is still weak in understanding the transition of camera viewpoints between two condition images. The output video looks like a transition of two slides if there is a big viewpoint change. Second, our method is affected by some flaws in the pretrained foundation models (e.g., watermark) because of the training-free nature. The upperbound of our method is constrained by the T2V model we used. Third, FlexTI2V is able to make the condition images as exact video frames at expected positions. Our method can not synthesize videos following only the subjects or background in the condition images (e.g., personalized video generation).
@@ D Limitation and Future Work
E Code and Data Release
We will publicly release our code, selected test data and textual prompts to facilitate future research in similar directions.

### v3

@@ B.6 Failure Cases
We demonstrate some failure cases of our method in Fig. C to show the boundary of the capability. We find it is challenging for our method to synthesize smooth camera viewpoint change in output videos (first example). If there is a big difference in viewpoints of two condition images, our method is very likely to treat the two images as irrelevant images and generates a hard transition without a reasonable motion. In addition, we also observe watermarks in a few output videos (second example). This is an inherent bias of the foundation model, which is inherited by our method. Hence, the performance of our method is still limited by the pretrained T2V foundation model.
Appendix C Implementation Details
Class Prompt
ApplyEyeMakeup “A person is applying eye makeup”
ApplyLipstick “A person is applying lip stick”
BabyCrawling “A baby is crawling”
BenchPress “A person is doing a bench press exercise”
Biking “A person is riding a bicycle”
BlowDryHair “A person is blow drying their hair”
BodyWeightSquats “A person is performing bodyweight squats”
BoxingPunchingBag “A person is punching a boxing bag”
BrushingTeeth “A person is brushing their teeth”
CuttingInKitchen “A person is cutting food in the kitchen”
Drumming “A person is playing drums”
Fencing “Two people are participating in fencing”
Haircut “A person is giving someone a haircut”
HandstandPushups “A person is performing handstand pushups”
HeadMassage “A person is doing a head massage for another one”
HorseRace “People are participating in a horse race”
HorseRiding “A person is riding a horse”
JumpRope “A person is jumping rope”
Kayaking “A person is kayaking in water”
Knitting “A person is knitting”
Mixing “A person is mixing ingredients”
MoppingFloor “A person is mopping a floor”
PlayingCello “A person is playing a cello”
PlayingFlute “A person is playing a flute”
PlayingGuitar “A person is playing a guitar”
PlayingPiano “A person is playing a piano”
PlayingViolin “A person is playing a violin”
PullUps “A person is doing pull-ups”
PushUps “A person is performing push-ups”
Rafting “People are rafting in a river”
RockClimbingIndoor “A person is climbing an indoor rock wall”
Rowing “People are rowing a boat”
ShavingBeard “A person is shaving their beard”
SkateBoarding “A person is skateboarding”
Skiing “A person is skiing down a slope”
Skijet “A person is riding a jet ski”
SkyDiving “A person is skydiving”
Surfing “A person is surfing on a wave”
TrampolineJumping “A person is jumping on a trampoline”
Typing “A person is typing on a keyboard”
WalkingWithDog “A person is walking a dog”
WritingOnBoard “A person is writing on a board”
@@ Appendix C Implementation Details
Table B: Prompt for each selected action class in UCF-101. More details of the test set is elaborated in Sec. C.1.
@@ C.1 Establishment of Test Set
Ni et al. [ni2024ti2v] select 10 classes in UCF-101 which are not diverse enough for a reliable evaluation. The textual prompts used in their work are also oversimplified. In this work, we further extend the test set to 42 classes with manually annotated textual prompt for each action. For each class, there are videos of 25 subjects performing the same action. For diversity, we randomly select one video of each subject in each action class, totaling 1050 test data samples. The prompts are listed in Tab. B.
@@ C.3 Implementation of Previous Methods
Appendix D Limitation and Future Work
Though our approach achieves great performance in many experimental settings compared with previous methods, we still observe some limitations as illustrated in Fig. C. First of all, FlexTI2V is still weak in understanding the transition of camera viewpoints between two condition images. The output video looks like a transition of two slides if there is a big viewpoint change. Second, our method is affected by some flaws in the pretrained foundation models (e.g., watermark) because of the training-free nature. The upperbound of our method is constrained by the T2V model we used. Third, FlexTI2V is able to make the condition images as exact video frames at expected positions. Our method can not synthesize videos following only the subjects or background in the condition images (e.g., personalized video generation).